Can you simplify the following expressions -
(square root of 36 t cubed) over 2t to the power of -2
And s to the power of 4 over the cubed root of 8s

Answer by stanbon(75887)

Can you simplify the following expressions -
(square root of 36 t cubed) over 2t to the power of -2
--------
sqrt[36t^3]/[2t^-2]
--------
= [6t*sqrt(t)]/[2/t^2]
--------
= [6t^3*sqrt(t)]/2
-----
= 3t^3*sqrt(t)
----
= 3t^(3+(1/2))
----
= 3t^(7/2)
=================
And s to the power of 4 over the cubed root of 8s
---
s^4/cbrt(8s)
-----
= s^4/(2*cbrt(s)
-----
= s^(4-(1/3))/2
-----
Cheers,
Stan H.
---------------
= (1/2)*s^(11/3)
============
Cheers,
Stan H.